# Identifying Network Structure of Linear Dynamical Systems: Observability and Edge Misclassification

### v1

@@ Abstract
This work studies the limitations of uniquely identifying a linear network’s topology from partial measurements of its nodes. We show that the set of networks that are consistent with the measurements are related through the nullspace of the observability matrix for the true network. In doing so, we illustrate how potentially many networks are fully consistent with the measurements despite having topologies that are structurally inconsistent with each other, an often neglected consideration in the design of topology inference methods. We then provide an aggregate characterization of the space of possible networks by analytically solving for the most structurally dissimilar network. We find that when observing over 6% of nodes in random network models (e.g., Erdős-Rényi and Watts-Strogatz) the rate of edge misclassification drops to $\sim$ 1%. Extending this discussion, we construct a family of networks that keep measurements $\epsilon$-“close” to each other, and connect the identifiability of these networks to the spectral properties of an augmented observability Gramian.
@@ I Introduction and Motivation
Delineating the causal relationships or connections in networks such as in groups of neurons or social interactions is of significant interest particularly when these connections are inferred from time series measurements of the nodes in the networks [1, 2, 3, 4, 5]. Despite efforts to develop different inference methods to accomplish this task, methods typically yield varying predictions [1]. Specifically, given some time series measurements of the nodal states, each method will infer a network topology that is consistent with the measurements, but often are structurally inconsistent with each other. Moreover, these methods typically assume that every node in the network can be measured and perturbed, an assumption that is often violated in practice for large scale networks where only a subset of the nodes are accessible. The motivation for this work spurs from interests arising in the field of neuroscience to understand the brain’s network structure referred to as the “connectome”, i.e., the connections between neurons or groups of neurons, and how this structure influences its function. Identifying the neural circuits can prove challenging as many variations to these circuits can display the same behavior [6, 7]. Existing systems theory literature studying parameter identifiability typically employs an input-output view [8, 9]. Furthermore, these notions of identifiability have been particularized to linear network models in which the transfer function of these models are studied [10], which has led to frameworks to understand how partially observed nodes impact identifiability of these systems [11, 12]. However, stimulating and measuring all neurons is infeasible and emphasizes the necessity for understanding limitations of partial measurements in a passive measurement setting, a gap this work aims to fill [^1] For results on network topology identification in nonlinear setting, see companion paper [13]..
@@
Figure 1: Given time series measurements of a selection of the nodes in a network, inference methods often predict different network topologies. This paper focuses on describing the set of possible networks consistent with these measurements.
This prompts a few pertinent questions: given a set of time series measurements, can we determine what other networks could have generated these measurements? Moreover, what is the most dissimilar network that could have generated these measurements? The focus of this work is to explore these questions in detail which in turn allows us to better understand the space of possible networks that could be inferred, we summarize this goal in Fig. 1.
@@
This work establishes necessary and sufficient conditions under which two systems will generate identical measurements when unperturbed by external inputs (Section II). In addition, we provide an aggregate characterization of the set of all networks that generate identical measurements, namely the most structurally dissimilar network (i.e., a proxy for how far networks can be from one another) and provide conditions under which this network can be analytically computed (Section III). We extend our analysis to the case where measurements are close but not necessarily identical establishing links to the spectral properties of an augmented observability Gramian (Section IV).
@@
Notation.  We denote matrices ($A$) and vectors ($x$) with uppercase and lowercase letters respectively. $A_{ij}$ denotes the element in the $i$’th row and $j$’th column of $A$. We denote the $\ell_{p}$ norm of a vector with $\|\cdot\|_{p}$. Moreover, we denote the element-wise (Hadamard) and Kronecker product of two matrices by $\odot$ and $\otimes$ respectively. Let $|\cdot|$ and $\text{sgn}(\cdot)$ be the absolute value and sign function respectively, which are applied element-wise. Let $\mathcal{N}(\cdot)$ and $\mathcal{R}(\cdot)$ correspond to the nullspace and range of a matrix. Lastly, let $\text{diag}(\cdot)$ be a diagonal matrix, and $\text{blkdiag}(\cdot)$ be a block diagonal matrix formed from the elements of their arguments.
@@ II Conditions for Identical Measurements
Consider a system that describes the linear dynamics of a network given as
$\Sigma:\begin{cases}\dot{x}(t)=Ax(t)&x(0)=x_{0}\\ y(t)=Cx(t).\end{cases}$ (1)
Here $A\in\mathbb{R}^{n\times n}$ denotes the system matrix and encodes the weighted adjacency matrix of the network. $x\in\mathbb{R}^{n}$ encodes the states of the nodes, and $y\in\mathbb{R}^{p}$ is a linear measurement of the states through an observation matrix $C\in\mathbb{R}^{p\times n}$. We study the conditions under which a perturbed system,
$\tilde{\Sigma}:\begin{cases}\dot{\tilde{x}}(t)=(A+\Delta)\tilde{x}(t)&\tilde{x}(0)=x_{0}\\ \tilde{y}(t)=C\tilde{x}(t),\end{cases}$ (2)
namely one that evolves under $A+\Delta$, where $\Delta$ is a perturbation to the network edge weights, yields identical measurements to the original system. We now make this notion precise.
@@
We say that the measurements of the original system (1) and the perturbed system (2) are indistinguishable if $\forall x_{0}\in\mathbb{R}^{n}$ $y(t)=\tilde{y}(t)~~\forall t\geq 0$.
@@
Measurements of the systems (1) and (2) are indistinguishable, if and only if $CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}$.
@@
Let $y(t)$ and $\tilde{y}(t)$ denote the measurements of (1) and (2) respectively. Assume that the measurements are indistinguishable, namely, $y(t)=\tilde{y}(t)~~\forall t$. Then we know that $y(t)=Ce^{At}x_{0}=Ce^{(A+\Delta)t}x_{0}=\tilde{y}(t)$. Moreover,
@@
Since $x_{0}$ is an arbitrary initial condition this is equivalent to requiring
@@
Since this equation holds $\forall t$, for necessity $(\!\!\implies\!\!)$, we examine that at time $t=0$ we must have that the derivatives and function values must be the same
@@
$\iff CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}.$
[…]
@@
and thus $CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}\implies y(t)=\tilde{y}(t)~~\forall t$, which concludes the proof. ∎
Immediately we can establish the following corollary.
@@
This corollary is a direct consequence of the Cayley-Hamilton Theorem. Namely, from Lemma 1, $y(t)=\tilde{y}(t)~~\forall t\iff CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}$. Notice that if $CA^{k}=C(A+\Delta)^{k}\text{ for }k=0,\cdots,n$ we have
$\begin{aligned} CA&=C(A+\Delta)\implies C\Delta=0\\ CA^{2}&=CA(A+\Delta)\implies CA\Delta=0\\ &~~\vdots\\ CA^{n}&=CA^{n-1}(A+\Delta)\implies CA^{n-1}\Delta=0.\end{aligned}$
Which is succinctly written as $CA^{k}\Delta=0$ for $k=0,\cdots,n-1$. Applying the Cayley-Hamilton theorem to express any higher powers of $A$ as a linear combination of $A^{k}$ with $k=0,\cdots,n-1$ by induction we may conclude $CA^{k}\Delta=0\quad\forall k\geq 0$. ∎
The result of Lemma 1 together with Corollary 1 can be rewritten in terms of the observability matrix $\mathcal{O}$ of the original system (1) $\Sigma$.
@@
From Corollary 1 we know that, $y(t)=\tilde{y}(t)~~\forall t\iff CA^{k}=C(A+\Delta)^{k}~~\text{ for }k=0,\cdots,n$. Moreover, for $k=1$ we see that this condition is satisfied if $C\Delta=0$, for $k=2$ we require $CA\Delta=0$, repeated evaluation establishes
$\begin{bmatrix}C^{\top}&(CA)^{\top}&\dots&(CA^{n-1})^{\top}\end{bmatrix}^{\top}\Delta=\mathcal{O}\Delta=0.$ (7)
@@
II-A Set of Possible Networks
Theorem 1 establishes that networks in the set, $\mathcal{A}=\{A+\Delta:\mathcal{O}\Delta=0\}$ all generate the same measurements. Now we consider a simple example to examine the cardinality of this set, consider the following network topology and measurement scheme
@@ II-A Set of Possible Networks
We note that the number of structurally dissimilar networks, i.e., networks that do not share the same underlying unweighted adjacency matrix, is $2^{n^{2}}$. For a network with $n=4$ nodes as in (8) there are $65536$ possible structurally dissimilar networks. We see that for the choice of $A$ and $C$ in (8)
@@
Hence, if we restrict our view to the first column of $A$, we can modify this column to be any of the following structurally dissimilar columns
@@
Notice, the edge $A_{11}=1$ is present in all topologies in $\mathcal{A}$, in fact, all of the incoming edges into node 1 ($A_{1j}~\forall j$) cannot be altered, this motivates formalizing this notion.
@@
Let the columns of $\Phi$ form a basis for $\mathcal{N}(\mathcal{O})$ and $e_{i}$ be a standard basis vector with a $1$ in the $i$’th component. We call an the edge encoded in $A_{ij}$ structurally essential if
@@
Notice that all of the edges encoded in $A_{1j}$ are structurally essential and cannot be altered. In fact, the entire row which corresponds to the evolution of node 1 is structurally essential.
@@
We call an edge encoded in $A_{ij}$ structurally decoupled from all other edges if $\exists v$ such that
@@
Moreover, we refer to edges that are not structurally decoupled as being structurally coupled to some other edges.
Notice, the edge encoded in $A_{41}$ was structurally decoupled from all other edges and was not structurally essential, hence we could freely remove or add this edge. The challenge in enumerating through all possible networks amounts to whether the basis of $\mathcal{N}(\mathcal{O})$ allows the structurally coupled edges in the network to be removed independently of one another, for instance, some structurally coupled edges may be only removed or added simultaneously. An explicit example is found in the fourth column of $A$, in this case the edges encoded in $A_{24}$ and $A_{34}$ are either both present or missing. Similarly, (10) shows that there is no way to simultaneously remove both $A_{21}$ and $A_{31}$. In this particular example there are $6^{3}\times 4=864$ structurally dissimilar networks in $\mathcal{A}$, emphasizing the unwieldy nature of enumerating this set. The combinatorial nature of iterating through changes to each edge motivates developing aggregate characterizations of the set $\mathcal{A}$. As a step towards doing so, we now consider a “worst-case” network that is useful in determining how dissimilar network topologies may be.
@@ III Dissimilar Networks
Having established Theorem 1, and given some true network $A$, we would like to understand the structure of the most dissimilar network that is consistent with the measurements made. Towards understanding this worst case network we formulate a minimization program over the nullspace of the observability matrix, specifically, we consider the following optimization program
@@
Where the matrix $Z\in\mathbb{R}^{n\times n}$ is defined such that $Z_{ij}=1$ if $A_{ij}\neq 0$ and $Z_{ij}=0$ otherwise. By defining $Z$ in this way and applying the element-wise product of $Z$ and $A+\Delta$ we are encouraging that the solution to this optimization problem will encourage removal of edges that exist in $A$. Using an $\ell_{1}$ norm allows us to encourage sparsity effectively setting these edges to zero.
@@
where $a=\text{vec}(A)\in\mathbb{R}^{n^{2}}$, $v=\text{vec}(V)\in\mathbb{R}^{n(n-r)}$ and $z=\text{vec}(Z)\in\mathbb{R}^{n^{2}}$, we can rewrite (13) as the following equivalent linear program (LP) [14]
$\min_{t,v}$ $\mathbbm{1}^{\top}t$ (16)
@@
We denote $\Gamma\coloneq\text{diag}(z)(I_{n}\otimes\Phi)\in\mathbb{R}^{n^{2}\times n(n-r)}$ for brevity.
Remark 2.
The constraints in (16) can be written as, $t\geq|a+\Gamma v|$, and the optimal solution $(t^{*},v^{*})$ satisfies $t^{*}=|a+\Gamma v^{*}|$. If it did not, we know $\exists s\geq 0~\text{s.t.}~t^{*}-s=|a+\Gamma v^{*}|$, but, $\mathbbm{1}^{\top}(t^{*}-s)<\mathbbm{1}^{\top}t^{*}$ and hence $t^{*}$ would not be optimal.
This program, (16), can be readily solved with LP solvers, however, we now provide conditions under which this problem admits a closed form analytic solution which is advantageous for large scale networks. We have found that it is typical for the solution of (13) to be identical to that of a relaxed $\ell_{2}$ program.

### v2

@@ Abstract
This work studies the limitations of uniquely identifying the structure (i.e., topology) of a networked linear system from partial measurements of its nodal dynamics. In general, many networks can be consistent with these measurements; this is a consideration often neglected by standard network inference methods. We show that the space of these networks are related through the nullspace of the observability matrix for the true network. We establish relevant metrics to investigate this space, including an analytic characterization of the most structurally dissimilar network that can be inferred, as well as the possibility of mis-inferring presence or absence of edges. In simulations, we find that when observing over 6% of nodes in random network models (e.g., Erdős-Rényi and Watts-Strogatz), approximately 99% of edges are correctly classified. Extending this discussion, we construct a family of networks that keep measurements $\epsilon$-close to each other, and connect the identifiability of these networks to the spectral properties of an augmented observability Gramian.
@@ I Introduction and Motivation
Inferring causal relationships between nodes in networks (e.g., regions in the brain) from time series measurements is broadly relevant to the engineering and scientific community [17, 4, 15, 19, 6]. Various inference methods have been proposed to accomplish this task, but they typically yield varying predictions [17]. Specifically, given some time series measurements of the nodal dynamics, different methods will infer different network structures (i.e., topologies); these are consistent with the measurements but often different from each other. Moreover, these methods typically assume that every node in the network can be measured and perturbed, an assumption that is often violated in practice for large networks where only a subset of nodes are accessible.
This question is particularly relevant for the field of neuroscience, where there is a concerted effort to understand the brain’s network structure (dubbed the connectome) based on time series measurements of brain regions and neurons/neural populations. A key question in the field is how the network structure of the brain influences its function. However, identifying network structure can prove challenging, as many different structures can exhibit similar behaviors [12, 14]. This challenge can be partially addressed by systems theory literature. The study of parameter identifiability typically employs an input-output view [10, 7]; these have been particularized to linear network models [18], which has led to frameworks to understand how partially observed nodes impact identifiability of these systems [11, 3]. However, existing work typically makes engineering-centric assumptions, e.g., assuming that we can excite all or most nodes with known perturbations. This is clearly infeasible in the neuroscience setting; most of the time even single perturbations are infeasible or prohibitively costly. The focus of this paper is to study the problem of network structure identification for linear dynamical systems under passive partial measurements. [^1] For results on network structure identification in the nonlinear setting, see companion paper [9]..
@@
Figure 1: Given time series measurements of a subset of nodes in a network, inference methods often predict different networks and structures. This paper describes the set of possible networks consistent with these measurements and provides characterizations of this set.
@@
In general, as implied above, many networks can be consistent with partial measurements. We are interested in identifying all such networks from given measurements (see Fig. 1). We start by establishing necessary and sufficient conditions under which two systems will generate identical measurements when unperturbed by external inputs (Section II). We also provide a formulation to solve for the most structurally dissimilar network (i.e., how bad can our inference be) and provide conditions under which this network can be analytically computed (Section III). We extend our analysis to the case where measurements are close but not necessarily identical, establishing links to the spectral properties of an augmented observability Gramian (Section IV).
@@
Notation.  $A_{ij}$ denotes the element in the $i$-th row and $j$-th column of $A$. The $\ell_{p}$ norm of a vector is denoted by $\|\cdot\|_{p}$. The element-wise (Hadamard) and Kronecker product by $\odot$ and $\otimes$. Let $|\cdot|$ and $\text{sgn}(\cdot)$ be the absolute value and sign function which are applied element-wise. Let $\mathcal{N}(\cdot)$ and $\mathcal{R}(\cdot)$ correspond to the nullspace and range of a matrix. Let $\text{diag}(\cdot)$ $\big(\text{blkdiag}(\cdot)\big)$ be a diagonal (block diagonal) matrix formed from the elements of its argument.
@@ II Conditions for Identical Measurements
Consider the following networked linear system:
$\Sigma:\begin{cases}\dot{x}(t)=Ax(t),&x(0)=x_{0}\\ y(t)=Cx(t).\end{cases}$ (1)
Where $A\in\mathbb{R}^{n\times n}$ encodes the weighted adjacency matrix of the network, $x\in\mathbb{R}^{n}$ encodes the states of the nodes, and $y\in\mathbb{R}^{p}$ is a linear measurement of the states through an observation matrix $C\in\mathbb{R}^{p\times n}$. We study the conditions under which a perturbed system,
$\tilde{\Sigma}:\begin{cases}\dot{\tilde{x}}(t)=(A+\Delta)\tilde{x}(t),&\tilde{x}(0)=x_{0}\\ \tilde{y}(t)=C\tilde{x}(t),\end{cases}$ (2)
yields identical measurements to the original system. Here, the perturbed system evolves under $A+\Delta$, where $\Delta$ is a perturbation to the network edges.
@@
The measurements of the original system (1) and the perturbed system (2) are indistinguishable if $\forall x_{0}$, $y(t)=\tilde{y}(t)~~\forall t\geq 0$.
@@
Measurements of systems (1) and (2) are indistinguishable, if and only if $CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}$.
@@
Let $y(t)$ and $\tilde{y}(t)$ denote the measurements of (1) and (2) respectively. Assume that the measurements are indistinguishable, i.e., $y(t)=\tilde{y}(t)~~\forall t$. Then,
@@
Since $x_{0}$ is an arbitrary initial condition this is equivalent to
@@
Since this equation holds $\forall t$, for necessity $(\Rightarrow)$, at time $t=0$ the derivatives and function values must be equal:
@@
$\Leftrightarrow CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}.$
Thus we have shown $y(t)=\tilde{y}(t)~~\forall t\Rightarrow CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}$. Now to establish sufficiency $(\Leftarrow)$, assume that $CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}$. Then we know that
@@
and thus $CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}\Rightarrow y(t)=\tilde{y}(t)~~\forall t$, which concludes the proof. ∎
The following corollary shows that we only require equivalence up to the $n$-th power.
@@
This corollary is a direct consequence of the Cayley-Hamilton Theorem. From Lemma 1, $y(t)=\tilde{y}(t)~~\forall t\Leftrightarrow CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}$. Notice that if $CA^{k}=C(A+\Delta)^{k}\text{ for }k=0,\cdots,n$ we have
$\begin{aligned} CA&=C(A+\Delta)\Rightarrow C\Delta=0\\ CA^{2}&=CA(A+\Delta)\Rightarrow CA\Delta=0\\ &~~\vdots\\ CA^{n}&=CA^{n-1}(A+\Delta)\Rightarrow CA^{n-1}\Delta=0.\end{aligned}$
Hence, $CA^{k}\Delta=0$ for $k=0,\cdots,n-1$. For all $k\geq n$, we can apply the Cayley-Hamilton theorem to express $CA^{k}\Delta$ as a linear combination of its lower powers, i.e., $CA^{k}\Delta=\sum_{m=0}^{n-1}\alpha_{m}CA^{m}\Delta$ for some $\alpha_{m}\in\mathbb{R}$. By the above analysis, all terms in this sum are equal to zero. Thus, we may conclude that $CA^{k}\Delta=0\quad\forall k\geq 0$. ∎
The result of Lemma 1 and Corollary 1 allow us to relate indistinguishability of (1) and (2) to observability of (1).
@@
From Corollary 1 we know that, $y(t)=\tilde{y}(t)~~\forall t\Leftrightarrow CA^{k}=C(A+\Delta)^{k}~~\text{ for }k=0,\cdots,n$. For $k=1$ we see that this condition is satisfied if $C\Delta=0$; for $k=2$ we require $CA\Delta=0$. Repeated evaluation establishes
$\begin{bmatrix}C^{\top}&(CA)^{\top}&\dots&(CA^{n-1})^{\top}\end{bmatrix}^{\top}\Delta=\mathcal{O}\Delta=0,$ (7)
@@
Theorem 1 implies the set of networks that are indistinguishable is $\mathcal{A}=\{A+\Delta:\mathcal{O}\Delta=0\}$. Consider the following example network and measurement scheme:
@@
The number of structurally dissimilar networks, i.e., networks that do not share the same underlying unweighted adjacency matrix, is $2^{n^{2}}$. For a network with $n=4$ nodes, as in (8), there are $65536$ possible structurally dissimilar networks. For the choice of $A$ and $C$ in (8)
@@
Consider the first column of $A$, which can be modified by adding a linear combination of vectors in $\mathcal{N}(\mathcal{O})$ into:
@@
Other values are possible; we have shown the five that represent all possible structures (i.e., sparsity patterns). Notice that edge $A_{11}=1$ is present in all structures in $\mathcal{A}$. In fact, all incoming edges into node 1 ($A_{1j}~\forall j$) cannot be altered, motivating a formalization of this phenomenon:
@@
Let the columns of $\Phi$ form a basis for $\mathcal{N}(\mathcal{O})$ and $e_{i}$ be a standard basis vector with $1$ in the $i$-th component. An edge encoded in $A_{ij}$ is structurally essential if
@@
In our example, all edges in $A_{1j}$ are structurally essential and cannot be altered without changing measured dynamics.
@@
An edge encoded in $A_{ij}$ is structurally decoupled from all other edges if $\exists v$ such that
@@
We refer to edges that are not structurally decoupled as being structurally coupled to some other edges.
Notice that edge $A_{41}$ is structurally decoupled from all other edges; thus, we can freely remove or add this edge. Here, enumerating all possible network structures amounts to determining whether the basis of $\mathcal{N}(\mathcal{O})$ allows the structurally coupled edges in the network to be removed independently of one another. Some structurally coupled edges may only be removed or added simultaneously. For example, edges $A_{24}$ and $A_{34}$ must both be present or both missing; similarly, (10) shows that there is no way to simultaneously remove both $A_{21}$ and $A_{31}$. In this particular example, there are $6^{3}\times 4=864$ structurally dissimilar networks in $\mathcal{A}$, emphasizing the unwieldy nature of enumerating this set. The combinatorial iterations through the various structures motivates developing aggregate characterizations of $\mathcal{A}$.
@@ III Dissimilar Networks
Given some true network $A$, we would like to understand the structure of the most dissimilar network that is consistent with the measurements made. Towards understanding this “worst-case” network we formulate a minimization program over the set $\mathcal{A}$. Defining $Z\in\mathbb{R}^{n\times n}$ such that $Z_{ij}=1$ if $A_{ij}\neq 0$ and $Z_{ij}=0$ otherwise, consider:
@@
The element-wise product of $Z$ and $A+\Delta$ alongside the use of an $\ell_{1}$ norm encourages the solution of (13) to remove edges that exist in $A$.
@@
where $a=\text{vec}(A)\in\mathbb{R}^{n^{2}}$, $v=\text{vec}(V)\in\mathbb{R}^{n(n-r)}$ and $z=\text{vec}(Z)\in\mathbb{R}^{n^{2}}$, we can rewrite (13) using the epigraph form of the problem to arrive at a linear program (LP) [5]:
$\min_{t,v}$ $\mathds{1}^{\top}t$ (16)
@@
Here $\Gamma=\text{diag}(z)(I_{n}\otimes\Phi)\in\mathbb{R}^{n^{2}\times n(n-r)}$.
Problem (16) can be readily solved with LP solvers. We now provide conditions under which this problem admits a closed form analytic solution; this provides better intuition about the problem and is also advantageous for the analysis of very large networks (which translate to very large LPs).